The following transactions were selected from among those completed by Bennett Retailers in November and December:

Nov. 20 Sold 20 items of merchandise to Customer B at an invoice price of $5,500 (total); terms 3/10, n/30.
25 Sold two items of merchandise to Customer C, who charged the $400 (total) sales price on her Visa credit card. Visa charges Bennett Retailers a 2 percent credit card fee.
28 Sold 10 identical items of merchandise to Customer D at an invoice price of $9,000 (total); terms 3/10, n/30.
29 Customer D returned one of the items purchased on the 28th; the item was defective and credit was given to the customer.
Dec. 6 Customer D paid the account balance in full.
20 Customer B paid in full for the invoice of November 20.

Required:

Assume that Sales Returns and Allowances, Sales Discounts, and Credit Card Discounts are treated as contra-revenues; compute net sales for the two months ended December 31. (Do not round your intermediate calculations.)

Net Sale=
Answer #1
Gross sales 14900 =5500+400+9000
Less: Sales returns -900
Less: Credit card fee -8 =400*2%
Less: Sales discount to Customer D -243 =(9000-900)*3%
Net Sales 13749